1) New cars are normal good and people’s incomes increase. Simultaneously, auto manufacturers must pay more for their worker’s health insurance. What is the effect on price and quantity of new cars?

 

2) The table below gives a supply schedule.

Point | Price (RM) | Quantity Supplied

A 5 10

B 15 30

C 25 50

D 35 90

 

a. Using the midpoint method, calculate the price elasticity of supply between points A and B; between B and C and between C and D.

b. Why is the elacticity of supply always positive?

c. Why does the elasticity of supply increase in value as more time passers after a price hike?


1
Expert's answer
2022-01-27T10:58:23-0500

a) An increase in income would lead to an increased demand for cars which are normal goods. This increased demand causes an increase in prices which is also led by increased pay for workers' health insurance which the suppliers charge to the buys in turn.

b) Midpoint Elasticity

A and B

Price Elasticity= % change in quantity demanded divided by % change in price

="\frac{(Q_B- Q_A)}{(Q_B+Q_A)/2}\times 100"


"\frac{(P_B- P_A)}{(P_B+P_A)/2}\times 100"


="\frac{(30- 10)}{(30+10)/2}\times 100=100" %


"\frac{(15- 5)}{(15+5)/2}\times 100=100"%


Price Ellasticity="\frac {100}{100}=1"

B and C

"\frac{(Q_C- Q_B)}{(Q_C+Q_B)/2}\times 100"


"\frac{(P_C- P_B)}{(P_C+P_B)/2}\times 100"


="\frac{(50- 30)}{(50+30)/2}\times 100=50" %


"\frac{(25- 15)}{(25+15)/2}\times 100=50"%

Price Elasticity= "\frac{50}{50}=" 1

C and D

"\frac{(Q_D- Q_B}{(Q_D+Q_B)2}"


"\frac{(P_D- P_C)}{(P_C+P_C)/2}\times 100"


"\frac{(90- 50)}{(90+50)/2}\times 100=\frac {400}{7}"%

"\frac{(35- 25)}{(35+25)/2}\times 100=\frac{100}{3}"


Price Elasticity= "(\frac{400}{7})\times (\frac{3}{100})=1.714"

b) Price elasticity is always positive because price and supply have a direct relationship. That is, they move in the same direction

c) With time, producers are able to adjust their production plans and hence the increased elasticity